gabethebabe:

Near San Fran, CA the vertical downward component of the earth's magnetic field is 4.8 x 10^-5 T. A car is traveling forwards at 25 m/s. An emf of 2.4 x 10^-3 V is induced between the sides of the car. a) which side of the car is positive b)what is the width of the car

gabethebabe:

I know this is a lenz law question but there is a hint that tells us to use epsilon= -N*(delta phi)/(delta Time)

gabethebabe:

B=4.8 x 10^-5 T V= 25m/s E= 2.4 x 10^-3 V

gabethebabe:

OH no im stupid this obviously requires us to use phi=B*A*cos(theta)

gabethebabe:

okay so using that we can say that phi is also = to B (L) (W) cos(theta) and we can assume cosine will be 0 for max emf

gabethebabe:

okay so we have a relation between the magnetic flux and the length but we cant solve that equation without relating it to either Velocity or the induced emf

gabethebabe:

this question would be so much easier if they gave me a function to derive freak me

cathrinerose23:

oka gabe slow down a sec

cathrinerose23:

the equations u gave earlier, are they the ones from the problem or the ones u came up with

gabethebabe:

they don't give me any equations except for that hint

gabethebabe:

but im deleting what ive ruled out

cathrinerose23:

cuz what im thinking is that the equations in the problem with the negative exponets we should try and solve them first to see if they take us somewhere closer

cathrinerose23:

dont delete anything yet

MrMudd183:

do what?

cathrinerose23:

solve this problem

MrMudd183:

i am so lost

cathrinerose23:

read the problem mudd